OpenStudy (spring98):

Part A: If (7^2)x = 1, what is the value of x? Explain your answer. Part B: If (7^0)x = 1, what are the possible values of x? Explain your answer.

OpenStudy (spring98):



OpenStudy (spring98):

@mathway @Nicki15 @Koikkara @logan13 @dm91 @AngelPSP

OpenStudy (anonymous):

for A, begin by solving 7^2, do you know how do do that/have access to a calculator?

OpenStudy (adilalvi):

Part A: If (7^2)x = 1 (7^2)x=1 (49)x=1 49x=1 x=1/49 x=0.0204082 answer Part B: If (7^0)x = 1 (7^0)x=1 (1)x=1 anything power 0 is equal to one. 1x=1 x=1 answer

OpenStudy (anonymous):

^^

OpenStudy (spring98):

49

OpenStudy (adilalvi):

what 49?

OpenStudy (spring98):

0.0204082 is the answer for part A?

OpenStudy (adilalvi):

yaah